what is the answer for this linear equation 10n+12=14n-12
Answer by drglass(89)

In a problem like this, you want to isolate the variable from the constants. This means, you want the n on one side and the numbers on the other. You can to this as follows:
subtract 10n from both sides to get
10n + 12 - 10n = 14n - 12 - 10n, rearrange the terms to get
10n - 10n + 12 = 14n - 10n - 12, or
0 + 12 = 4n - 12
Now, add 12 to both sides
12 + 12 = 4n - 12 + 12 or
24 = 4n
Divide both sides by 4 to get
24/4 = 4n/4 or
6 = n.
You can check this by replacing n with 6 in the original equation.
10(6) + 12 = 14(6) - 12
60 + 12 = 84 - 12
72 = 72
This is true, so n = 6 is correct